OpenStudy (anonymous):

need help solving this logarithmic e^4x-5-7=11243

OpenStudy (zepp):

What you want to do is to isolate the exponential variable, here I'll show you the process\[e^{4x}-5-7=11243 \\e^{4x}-12=11243 \\e^{4x}=11243+12 \\e^{4x}=11255 \]As you can see here, we get e to the something is equal to 11255, we can also say that the natural logarithm of 11255 is equal to that something \(\ln(11255)=4x\) And voila.